OpenStudy (anonymous):

4. Part A: Explain why the x-coordinates of the points where the graphs of the equations y=2-x and y = 4x + 3 intersect are the solutions of the equation 2-x = 4x + 3 Hint: Think about graphing (5 points) Part B: Make tables to find the solution to 2-x = 4x + 3. Take the values of x between −3 and 3. (5 points) @bruhhh

OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 2-x and y = 4x + 3 intersect are the solutions of the equation 2-x = 4x + 3. Hint: Think about graphing (5 points) Part B: Make tables to find the solution to 2-x = 4x + 3. Take the values of x between −3 and 3. Be sure to state the solution. (5 points) y = 2-x y = 4x + 3 x y 1 0 -1 -2 x y 1 0 -1 -2

OpenStudy (anonymous):